# A Survey of Efficient Reasoning for Large Reasoning Models: Language, Multimodality, and Beyond

### v1

@@ 7.4 Building Efficient Reasoning Applications
RAG Gao et al. (2023); Zhao et al. (2024b); Sun et al. (2024b); Qu et al. (2024a) offers a straightforward and effective approach to addressing the inherent limitations of static parameters in generative models by retrieving content from external knowledge bases Su et al. (2024c). Recently, agentic RAG systems have empowered models to autonomously determine when and what knowledge to retrieve, demonstrating enhanced planning and problem-solving capabilities Chen et al. (2024a); Li et al. (2025c). Li et al. (2025c) further combine large retrieval models (LRMs) with an agentic RAG mechanism, incorporating a “Reason-in-Documents” module to refine retrieved content. They enable dynamic retrieval of external knowledge when LRMs face uncertainty. Additionally, Wang et al. (2025a) introduce a method for training O1-like RAG models that perform step-by-step retrieval and reasoning over relevant information before generating the final answer. During inference, they modulate the model’s computational cost by adjusting the length and number of sampled retrieval chains. This integration of retrieved content enhances the depth and breadth of LRMs’ reasoning, highlighting the importance of efficient reasoning in RAG systems.
@@
In the agent scenarios, Zhou et al. (2025c) discover that LRMs outperform LLMs in reasoning-intensive tasks, such as Plan Design, by leveraging iterative reflection to achieve superior results. However, their reliance on extensive reasoning often incurs significant computational overhead, thus reducing efficiency in time-critical contexts Li et al. (2024c); Zhang et al. (2024b). Cuadron et al. (2025) further study the overthinking phenomenon in the magnetic tasks, and elevate overthinking scores correlate negatively with performance. They further propose to select solutions with lower overthinking scores to mitigate the overthinking and achieve superior results. Thus, achieving efficient reasoning for agents is a promising avenue for research.
@@
In addition to the aforementioned challenges, enhancing tool efficiency in LRMs necessitates a multifaceted approach. One promising direction is the incorporation of hierarchical reasoning and early exit strategies that dynamically terminate computations once sufficient confidence is achieved, thereby reducing unnecessary function calls Qin et al. (2024a). Furthermore, parallel execution schemes, which can be accelerated by specialized hardware such as GPUs or FPGAs, further mitigate latency Qu et al. (2025a). Another avenue involves dynamic query routing, where the system adjusts the complexity of its planning and reasoning processes based on current task demands and resource availability Gao et al. (2024). In summary, these integrated strategies form a coherent framework that optimizes both response speed and performance, paving the way for the efficient and practical deployment of LRMs in real-world applications Hadi et al. (2023).
@@
Considering the nascent state of efficient reasoning research in LRMs, numerous other compelling research directions warrant further exploration, such as efficient reasoning for coding Yang et al. (2025b); Jiang et al. (2024), autonomous driving Yurtsever et al. (2020); Zhao et al. (2024a), health care Temsah et al. (2024, 2025), and embodied AI Duan et al. (2022); Liu et al. (2024d).
@@ 7.5 Evaluation and Benchmark
Currently, most studies Aggarwal and Welleck (2025); Xia et al. (2025) evaluate the efficiency of Large Language Models (LRMs) on complex math problems, such as GSM8K Cobbe et al. (2021a), MATH Hendrycks et al. (2021), ASDIV Miao et al. (2021), and AIME. These works primarily focus on comparing the trade-offs between accuracy and token consumption across different methods.
To investigate the phenomenon of overthinking in math problems, where multiple solutions are generated for a single question, Chen et al. (2025c) analyze each solution and propose two efficiency metrics that address both outcome and process perspectives. These metrics are designed to assess the computational resource efficiency of o1-like LRMs. Specifically, the outcome efficiency metric evaluates how much subsequent solutions improve accuracy beyond the first solution. In contrast, the process metric assesses how much subsequent solutions contribute to the diversity of solutions, independent of correctness Zeng et al. (2024b); Song et al. (2025b). Future research could delve deeper into the analysis of the immediate solutions generated, particularly by examining how each initial response contributes to the overall efficiency of the reasoning process Bi et al. (2024).
In addition to straightforward math problems, it is crucial to explore whether efficient reasoning compromises the level of intelligence. Therefore, it is essential to evaluate efficient reasoning in more general domains, such as creativity and innovation Lu et al. (2024a); Ruan et al. (2024); Franceschelli and Musolesi (2024). Towards the evaluation of efficient reasoning of LRMs, Hashemi et al. (2025) introduce DNA Bench, a benchmark designed to expose a vulnerability in LRMs’ tendency for over-reasoning. The prompts in this benchmark are carefully crafted to mislead LRMs into generating excessively verbose reasoning chains, ultimately leading to incorrect responses. This highlights the need for more nuanced approaches to measure and improve reasoning efficiency in LRMs. Looking ahead, designing more comprehensive and diverse benchmarks for efficient reasoning presents a promising avenue for future research.

### v2

@@ 7.4 Building Efficient Reasoning Applications
RAG gao2023retrieval; zhao2024retrieval; sun2024surf; qu2024alleviating offers a straightforward and effective approach to addressing the inherent limitations of static parameters in generative models by retrieving content from external knowledge bases su2024conflictbank. Recently, agentic RAG systems have empowered models to autonomously determine when and what knowledge to retrieve, demonstrating enhanced planning and problem-solving capabilities chen2024mindsearch; li2025search. li2025search further combine large retrieval models (LRMs) with an agentic RAG mechanism, incorporating a “Reason-in-Documents” module to refine retrieved content. They enable dynamic retrieval of external knowledge when LRMs face uncertainty. Additionally, wang2025chain introduce a method for training O1-like RAG models that perform step-by-step retrieval and reasoning over relevant information before generating the final answer. During inference, they modulate the model’s computational cost by adjusting the length and number of sampled retrieval chains. This integration of retrieved content enhances the depth and breadth of LRMs’ reasoning, highlighting the importance of efficient reasoning in RAG systems.
@@
In the agent scenarios, zhou2025largereasoningmodelsagent discover that LRMs outperform LLMs in reasoning-intensive tasks, such as Plan Design, by leveraging iterative reflection to achieve superior results. However, their reliance on extensive reasoning often incurs significant computational overhead, thus reducing efficiency in time-critical contexts li2024personal; zhang2024towards. cuadron2025dangeroverthinkingexaminingreasoningaction further study the overthinking phenomenon in the magnetic tasks, and elevate overthinking scores correlate negatively with performance. They further propose to select solutions with lower overthinking scores to mitigate the overthinking and achieve superior results. Thus, achieving efficient reasoning for agents is a promising avenue for research.
@@
In addition to the aforementioned challenges, enhancing tool efficiency in LRMs necessitates a multifaceted approach. One promising direction is the incorporation of hierarchical reasoning and early exit strategies that dynamically terminate computations once sufficient confidence is achieved, thereby reducing unnecessary function calls qin2024toollearningfoundationmodels. Furthermore, parallel execution schemes, which can be accelerated by specialized hardware such as GPUs or FPGAs, further mitigate latency qu2025tool. Another avenue involves dynamic query routing, where the system adjusts the complexity of its planning and reasoning processes based on current task demands and resource availability gao2024efficient. In summary, these integrated strategies form a coherent framework that optimizes both response speed and performance, paving the way for the efficient and practical deployment of LRMs in real-world applications hadi2023survey.
@@
Considering the nascent state of efficient reasoning research in LRMs, numerous other compelling research directions warrant further exploration, such as efficient reasoning for coding yang2025code; jiang2024survey, autonomous driving yurtsever2020survey; zhao2024autonomous, health care temsah2024openai; temsah2025deepseek, and embodied AI duan2022survey; liu2024aligning.
@@ 7.5 Evaluation and Benchmark
Currently, most studies aggarwal2025l1controllinglongreasoning; xia2025tokenskip evaluate the efficiency of Large Language Models (LRMs) on complex math problems, such as GSM8K cobbe2021gsm8k, MATH hendrycksmath2021, ASDIV miao2021diverse, and AIME. These works primarily focus on comparing the trade-offs between accuracy and token consumption across different methods.
To investigate the phenomenon of overthinking in math problems, where multiple solutions are generated for a single question, chen2025think23overthinkingo1like analyze each solution and propose two efficiency metrics that address both outcome and process perspectives. These metrics are designed to assess the computational resource efficiency of o1-like LRMs. Specifically, the outcome efficiency metric evaluates how much subsequent solutions improve accuracy beyond the first solution. In contrast, the process metric assesses how much subsequent solutions contribute to the diversity of solutions, independent of correctness zeng2024mr; song2025prmbench. Future research could delve deeper into the analysis of the immediate solutions generated, particularly by examining how each initial response contributes to the overall efficiency of the reasoning process bi2024forest.
In addition to straightforward math problems, it is crucial to explore whether efficient reasoning compromises the level of intelligence. Therefore, it is essential to evaluate efficient reasoning in more general domains, such as creativity and innovation lu2024llm; ruan2024liveideabench; franceschelli2024creativity. Towards the evaluation of efficient reasoning of LRMs, hashemi2025dnabenchsilencesmarter introduce DNA Bench, a benchmark designed to expose a vulnerability in LRMs’ tendency for over-reasoning. The prompts in this benchmark are carefully crafted to mislead LRMs into generating excessively verbose reasoning chains, ultimately leading to incorrect responses. This highlights the need for more nuanced approaches to measure and improve reasoning efficiency in LRMs. Looking ahead, designing more comprehensive and diverse benchmarks for efficient reasoning presents a promising avenue for future research.